# Cocoa: Co-Planning and Co-Execution with AI Agents

### v2

@@ 7.1. Chat, Interactive Plans, or Both?
In practice, rather than choosing between either interactive plans or chat when building future systems, we see research opportunities for AI agent interfaces to strategically combine the two and harness the best of both worlds, especially when assisting the user with complex, long-running requests. For example, in addition to using chat and interactive plans sequentially, we can also consider how they can be integrated within one another. Embedding chat within interactive plans can provide more flexibility and organization. For example, during co-planning, chat-based features can be used to specify higher-level desiderata for the proposed plans when none are satisfactory for the user. Additionally, when the agent re-plans, users may guide the re-planning process via chat to ensure alignment with their objectives. During co-execution, the user may engage in conversation with the agent at every step to iterate on outputs. While direct manipulation of step outputs was desirable to our participants, some also did not appreciate the agent overwriting old outputs as the step was rerun. A conversation thread for each step can maintain iteration history while allowing users to also specify output modifications or transformations in natural language. This way, the interactive plan also serves as a way to organize conversation threads with the agent as the agent operates over many plan steps.
Conversely, interactive plans can be incorporated into chat-first interfaces to provide more structure and steerability for long-running tasks. ChatGPT, when powered with OpenAI’s o1 model, offers expandable chain-of-thought activity traces as the agent reasons at inference time within a chat environment (OpenAI, 2024). While these activity traces are not yet interactive, interactivity can be valuable when the agent is operating in information-poor environments where the user has more expertise and context than the agent. With interactive plans, the agent exposes critical points of operation that can benefit from user steering, such that the agent can continue with higher confidence that it is not working with erroneous information. After the agent produces a final output, the user can collapse the plan and continue the conversation. Returning to a chat environment after interactive planning can be useful for situations in which the user is mostly interested in open-ended exploratory problems, with occasional periods of structured execution.
@@ 7.2. The Practical Significance of User Steps and Advanced Planning
In Cocoa, users are presented with a multi-step interactive plan up front, which the agent adheres to and can modify while operating. This plan also allows users to delegate a step to themselves or the agent. Both design decisions are a departure from many existing agentic systems, where the agent dynamically constructs a plan step-by-step as it generates outputs (Yao et al., 2022b) and seeks user input only when it is unable to complete a task (Wu, 2024). We discuss two practical reasons—cost and safety—for why such a departure may be desirable when using agents in the real world.
@@
AI agents can be prohibitively expensive to run (Kapoor et al., 2024; Yang et al., 2024). Because an agent’s actions often involve multiple LLM calls, costs can quickly accumulate, especially for more complex and long-running tasks. In response, academic projects have implemented cost-capping measures to alleviate this (Yang et al., 2024), and researchers have called for agent evaluations to be cost-controlled (Kapoor et al., 2024). In user-facing and interactive systems, cost also includes time spent waiting for agents to come back with results. These cost considerations can also change how the user interacts with the agent. When participants made the decision to assign most or all steps to the agent, they viewed the agent’s actions as rather inconsequential and low-cost—they could run the agent and see what happens, and edit the output post hoc (Section 6.3.1). However, while this was the case in user studies, it may not be true in many real-world contexts where users and/or developers may need to pay monetary and time costs for each agent step. In this case, they may see user steps as an important cost-saving measure, especially if the step is not particularly burdensome for the user, but may be difficult for current AI agents to complete effectively. Efficiently leveraging human effort and expertise where appropriate via user steps can be a key step towards developing cost-aware agents that are usable and affordable in practice.
@@
The increased autonomy of agentic systems also comes with increased safety concerns, as AI harms become more difficult to anticipate and accountability for AI actions become harder to trace (Chan et al., 2023, 2024). Interactive plans not only provide more transparency into agent actions (Section 6.1.3), but they provide some assurance about and control over the agent’s execution trajectory. By generating a plan ahead of execution that the agent adheres to, problems in AI safety such as harm anticipation become much more tractable. While step-by-step plan generation used by existing agent frameworks such as ReAct (Yao et al., 2022b) may also accomplish a similar goal by asking for user approval whenever a new step is generated, this can severely limit agent autonomy when it is desirable and safe, and posing unnecessary supervisory burden onto the user. Additionally, advanced planning enables assessments of task risk pre-execution. Plan steps determined to be of higher risk (e.g., requires working with passwords) can be assigned to the user—even if the agent is capable of completing them—as a risk mitigation measure.
7.3. Interactive Plans Outside of Documents
From our user study, we learned that participants found interactive plans to be especially valuable for working with AI agents to tackle partially- or well-defined problems with a structured, organized approach (Section 6.2). Cocoa supports these approaches through co-planning and co-execution in a document environment, but there are many other environments and user groups beyond researchers that may benefit from the same interaction design pattern. Here, we paint our visions for how interactive plans may effectively interleave human and AI agency in three non-document contexts: workplace messaging apps (e.g., Slack), software development environments (e.g., code editors), and UI design tools (e.g., Figma).
In a Slack channel, interactive plans help a team coordinate efforts between channel members, including an AI agent, when working towards a common goal in a complex project. For example, when the project lead proposes starting a new phase in the project, they can invoke the AI agent to initiate co-planning. The AI agent uses the channel’s conversation history and media to generate an initial plan of action and present this plan as a widget, similar to Slack widgets for interactive polling. Team members can edit descriptions of the plan steps and assign themselves, each other, and/or the agent to each step. The team works through the plan step by step, with the plan offering a clear visualization of progress and blockers. Clicking into each step allows members to access step-specific artifacts, discussion threads, and media created by those assigned to that step. As the plan may easily shift with progress, the agent evaluates the remainder of the plan at the completion of each step and proposes changes to the plan, with proposals grounded in team outputs from previous steps. The interactive plan becomes a centralized, living artifact that a team can use flexibly orchestrate their work.
In an AI-powered code editor, an interactive plan can help efficiently harmonize coding efforts of humans and AI. Programming itself involves series of a highly structured and logical activities, and are thus well-suited for interactive plans. A programmer can express their intent for implementing a new feature that involves code changes across multiple files, and the agent first proposes an interactive plan for doing so. The programmer can review this plan and assign some tasks for the agent and some for themselves. This is crucial for both efficiency and reliability. While recent advances in autonomous software engineering agents (e.g., Devin (Wu, 2024)) exhibit impressive capabilities to write and edit code, they can get stuck on or are slow to perform simple tasks a human software engineer can easily and quickly accomplish, such as minor UI styling adjustments (Emanuilov, 2024). Moreover, these agents can be costly to operate, both in terms of tokens and time, potentially spending hours on a task. By strategically delegating plan steps, the developer can cover for the agent’s limitations (and vice versa), and parallelize their development efforts with those of the agent on long-running tasks. The agent’s code can also be unreliable and non-functional. The developer may self-assign some steps that involve code review or writing tests to ensure the agent is not jeopardizing the reliability of the codebase.
In a Figma canvas, interactive plans can help interface designers work with a multimodel AI agent to collaboratively design UI mockups and prototypes. The designer may already have some design specifications and low-fidelity wireframes on their canvas. To create a set of more comprehensive and higher-fidelity mockups, the designer can invoke the agent and edit an interactive plan that the agent produced using canvas content as context. Upon inspecting the plan, the designer may realize the need to specify the exact wireframes for the agent to work with due to the clutter of the canvas, and create an additional user step accordingly. The designer may also have an existing design system they would like the agent to use when transform the wireframes into high-fidelity mockups, so they create another user step in which they link to their design system. As the agent completes its steps and deposits new designs into the canvas, the designer can edit the agent’s work as they see fit. The agent then automatically completes the otherwise menial and laborious task of connecting the mockups together into an interactive prototype. The designer is mostly satisfied with the results, but wants the agent to replicate a similar interaction from one of their past designs. They add a user step where they link to that design and rerun the agent’s prototyping step. Overall, interactive plans provide a streamlined workflow for the designer to guide and collaborate with an AI agent in their design canvas.
8. Limitations
While we controlled our user study to be 90 minutes long to avoid fatigue, some literature-grounded tasks may still require more time to complete. The time-constrained nature of our study impacted how participants interacted with our system, especially their decision of whether to delegate plan steps to the agent (Section 6.3.1). In the near future, we hope to run a longitudinal deployment of Cocoa to study the use of interactive plans in-the-wild as researchers use them to work on in-progress projects. Related to step assignment, the default assignment of user and agent steps in Cocoa is a result of researchers’ preferences from our formative studies. Past work explored automated methods for a model to decide when to defer a task to a human expert versus acting on its own (Mozannar and Sontag, 2020). Future work may investigate applications of these methods in interactive plans.
@@ 8. Limitations
Based on our formative study, we built Cocoa as a document editor because documents are natural sites of planning for researchers and present an ideal environment for agent interaction (Section 3.3.5). However, the linear nature of a document also has its limitations. Some participants wished to “fork” plans and interactively run multiple versions of a plan in parallel. These types of interactions are better supported by node-based canvases, which have been gaining popularity as a means of sensemaking and creative exploration with LLMs (e.g., (Suh et al., 2023; Pu et al., 2024; Arawjo et al., 2023)). As briefly mentioned in Section 7.3, future work can explore embedding interactive plans in a canvas environment such as Figma.
@@
The underlying agent used in Cocoa also had some technical limitations. Since it was powered by a language-only model, it was not capable of taking visual content (e.g., figures or slides) as input. The agent was also not capable of executing code—a capability which a couple participants inquired about during the study. While it is sufficient for this current work to explore interactive plans as a novel interaction paradigm, future work can expand their utility with multimodal agents and/or agents that can execute code.
@@
Finally, our participants were researchers in CS and CS-adjacent areas. Research culture and incentives specific to CS may bias our results and limit our imagination of how interactive plans can be used in other areas. P10, who used to work in wet labs in their undergraduate research, shared that a useful application of interactive plans would be helping wet lab researchers walk through the steps necessary to prepare for lab experiments. Future work may investigate novel applications of interactive plans in domains beyond CS.
@@ 9. Conclusion
In this paper, we introduced Cocoa, an interactive document editing environment for scientific researchers to tackle open questions and tasks within their research projects with an AI agent. Cocoa introduced a new interaction pattern that enabled Co-planning and Co-execution interactive plans between a researcher and an AI agent. Our formative study with 9 researchers informed the design and behavior of Cocoa. After implementing Cocoa, we evaluated it through a user study with 16 researchers and found that when compared to a chat baseline powered by the same underlying AI agent. Results showed that Cocoa’s novel UI affordances allowed researchers to more flexibly steer the agent’s behavior in helpful directions and without increasing user effort (i.e., ease of use). We also learned, through our study’s qualitative data, researchers’ tasks-specific preferences for when they prefer interactive plans over chat. Informed by our results, we discuss numerous practical implications for the design and implementation of agentic AI systems, from approaches that combine the benefits of interactive plans and chat, to environments for interactive planning beyond documents. Overall, our contributions set the stage for new interactive paradigms that foster effective collaboration between humans and AI agents.
Agentic AI systems have exciting potential to transform our digital experiences and advance long-standing visions held by HCI and AI communities. However, these transformations and advancements also demand renewed attention to strategies for ensuring effective collaboration between human users and AI agents in the real world. Our work advances emerging efforts in this area.
@@ Appendix A Frequently Asked Questions
A.1. What’s the novelty of this work?
We introduce a new interaction design pattern—interactive plans—that allow humans and AI agents to work together on information-intensive tasks. Interactive plans support a human-AI workflow that interleaves co-planning (collaboratively drafting a plan to complete a task) and co-execution (collaboratively executing a drafted plan). We demonstrate, by implementing interactive plans into Cocoa, that our pattern can improve agent steerability for users while preserving ease of use, when compared to conventional chat-based interfaces for agents (see Section 6.1). Our work also contributes user feedback on when particular design patterns for agent interfaces may be (in)appropriate (see Section 6.2). Our work contributes to ongoing efforts to improve the usability of AI agents and ensure responsible, safe, and effective real-world deployment.
A.2. What are the main differences between interactive plans and chat interfaces?
Interactive plans allow users to work with the agent in much more direct ways (hence the improved steerability) compared to chat. Technically speaking, both interfaces allow the agent to start with a plan and allow the user to correct agent outputs. However, in a chat-based system, this can be cumbersome—users have to specify the changes they want to make and wait for the agent to make them (with success not guaranteed) instead of doing it themselves directly. Further, we recognize that plans can easily change upon seeing completed steps’ outputs. That’s why Cocoa supports automatic replanning (see Fig. 6) to dynamically modify the plan based on user edits. A helpful analogy here is computational notebooks. Interactive plans allow agent workflows to be crafted and executed interactively like cells of code in a notebook, rather than running the entire program at once like in chat.
Of course, that’s not to say that chat is an inferior interface—we learned about the pros and cons of both interfaces through our user studies and discuss how to combine the best of both worlds in Section 7.1.
A.3. Why is important to get the user involved? Can’t we just let AI automate everything?
@@
A.4. Why didn’t you use ChatGPT or a similar publicly available AI chatbot for your baseline?
Our baseline’s interface closely matches that of ChatGPT and other AI chatbots, but the underlying system is slightly different. We use an LLM agent powered by GPT-4o, but specifically trained it to use the Semantic Scholar API as tools to reliably perform research-related tasks such as searches for papers/authors/topics, sorting papers by various metrics including citation count, summarizing papers, comparing papers, retrieving prominent papers by a particular author, etc. See Section 4.4 for technical details. However, our main contribution is not building a new LLM agent nor advancing agent capabilities—rather, it is to explore cost and benefit tradeoffs between interaction techniques for humans to work with AI agents. We use the same underlying agent in Cocoa, so we kept this factor constant across our system and baseline.
A.5. How are the agent’s plans generated?
We conducted a formative study with 12 researchers where we interviewed them about their research processes and synthesized a repository of common steps. These steps were developed into higher level templates and used as few-shot learning examples for how to compose steps into a plan in the agent’s system prompt. See Section 3 for more details about our formative study and Appendix F for our prompts.

### v3

@@ 8.1. Chat, Interactive Plans, or Both?
In practice, rather than choosing between either interactive plans or chat when building future systems, we see research opportunities for AI agent interfaces to strategically combine the two, especially when assisting the user with complex, long-running requests.
Consider “deep research” systems that assemble a detailed report on a particular topic upon searching for and synthesizing information from around the web (Google, 2025; OpenAI, 2025a; Perplexity, 2025). Some of these systems may clarify user requests in a chat conversation (OpenAI, 2025a) or allow the user to prompt the agent to devise a different plan (Google, 2025) before starting the research process, but neglect user involvement thereafter. Integrating interactive plans can be particularly valuable here because they afford iteration and control over specific parts of planning and execution. For example, the user may not be satisfied with the sources the agent is using to inform its answers and wants to steer the agent to use information from a narrow set of websites for a particular plan step. This kind of interaction is expensive without interactive plans because it requires re-running the entire research process. Interactive plans allow the user to focus on improving outcomes of a particular step by editing that step and rerunning it or directly editing the agent’s outputs at that step. Rather than re-executing the entire plan, the agent updates only the subsequent plan steps with the new user input. After the interactive plan has finished executing, users may return to interacting with the agent in a chat interface to ask questions about specific outputs or steps. These are precisely the benefits that computational notebooks provide for data analysts (Choi et al., 2023; Ayobi et al., 2023; Chattopadhyay et al., 2020).
In addition to using chat and interactive plans sequentially, we can also consider how they can be integrated within one another. Embedding chat within interactive plans can provide more flexibility and organization. For example, during co-planning, chat-based features can be used to specify higher-level desiderata for the proposed plans when none are satisfactory for the user. During co-execution, the user may engage in conversation with the agent at every step to iterate on outputs. This way, the interactive plan also serves as a way to organize conversation threads with the agent as the agent operates over many plan steps. Conversely, interactive plans can be incorporated into chat-first interfaces to provide more transparency and steerability. This can be useful for open-ended exploratory problems, with occasional periods of structured execution.
@@ 8.2. The Practical Significance of User Steps and Advanced Planning
In Cocoa, users are presented with a multi-step interactive plan up front, which the agent adheres to and can modify while operating. This plan also allows users to delegate a step to themselves or the agent. Both design decisions are a departure from many existing agentic AI systems, where the agent dynamically constructs a plan step-by-step as it generates outputs (Yao et al., 2022b) and seeks user input only when it is unable to complete a task (Wu, 2024). We discuss two practical reasons—cost and oversight—for why such a departure may be desirable for practical agents in the real world.
@@
AI agents can be prohibitively expensive to run (Kapoor et al., 2024; Yang et al., 2024). Because an agent’s actions often involve multiple LLM calls, costs can quickly accumulate, especially for more complex and long-running tasks. In response, academic projects have implemented cost-capping measures to alleviate this (Yang et al., 2024), and researchers have called for agent evaluations to be cost-controlled (Kapoor et al., 2024). In user-facing interactive systems, cost also includes time spent waiting for agents to come back with results. These cost considerations change user behavior. When participants made the decision to assign most or all steps to the agent in our lab studies, they viewed the agent’s actions as rather inconsequential and low-cost—they could run the agent and see what happens, and edit the output post hoc (Section 6.3). However, while this was the case in the lab, it may not be true in some real-world contexts where users and/or developers need to pay monetary and time costs for each agent step. The user steps become an important cost-saving measure. User can choose to complete a step that is not particularly burdensome or is too difficult for current AI agents to complete effectively. Efficiently leveraging human effort and expertise where appropriate via user steps can be a key step towards developing cost-aware agents that are usable and affordable in practice.
@@
The increased autonomy of AI agents also comes with increased risks, as AI harms become more difficult to anticipate, and accountability for AI actions becomes harder to trace (Chan et al., 2023, 2024). Interactive plans not only provide more transparency into agent actions, but also provide some assurance about and control over the agent’s execution trajectory. By generating a plan ahead of execution that the agent adheres to, harm anticipation and mitigation become more tractable. While step-by-step plan generation used by existing agent frameworks, such as ReAct (Yao et al., 2022b), may also accomplish a similar goal by asking for user approval whenever a new step is generated, this can severely limit agent autonomy when desirable and may also disengage the user if their only form of supervision is repetitive rubber-stamping (Kapoor et al., 2024). Additionally, advanced planning enables assessments of task risk pre-execution. Plan steps determined to be of higher risk (e.g., requiring working with passwords) can be assigned to the user—even if the agent is capable of completing them—as a risk mitigation measure.
9. Limitations and Future Work
@@
Based on our formative study, we built Cocoa as a document editor because documents are natural sites of planning for researchers and present an ideal environment for agent interaction (Section 3.2.1). However, the linear nature of a document also has its limitations and does not support “forking” plans and iterating on versions of a plan in parallel. These types of interactions are better supported by node-based canvases, which have been gaining popularity as a means of sensemaking and creative exploration with LLMs (e.g., (Suh et al., 2023; Pu et al., 2024; Arawjo et al., 2023)). Future work can explore new interaction techniques for interactive plans. The default assignment of user and agent steps in Cocoa is also a result of researchers’ preferences from our formative study. Past work explored automated methods for a model to decide when to defer a task to a human expert versus acting on its own (Mozannar and Sontag, 2020). Future work may investigate applications of these methods in interactive plans and/or leverage inference-time scaling to improve plan quality (OpenAI, 2024; Muennighoff et al., 2025).
@@
The underlying agent used in Cocoa also had some technical limitations. Since it was not multimodal, it was not capable of taking visual content (e.g., figures or slides) as input. The agent was also not capable of executing code—a capability which a couple of participants inquired about during the study. While it is sufficient for this current work to explore interactive plans as a novel interaction paradigm, future work can expand their utility with multimodal agents and/or agents that can execute code.
@@
Another limitation is the demographics of our participants. Our participants were researchers in CS and CS-adjacent areas. Research culture and incentives specific to CS may bias our results and limit our imagination of how interactive plans can be used. P10, who used to work in wet labs in their undergraduate research, shared that a useful application of interactive plans would be helping wet lab researchers walk through the steps necessary to prepare for lab experiments. Future work may investigate novel applications of interactive plans in domains beyond CS.
While our work only explored the use of interactive plans by a single user, one interesting future direction is to explore interactive plans in multi-user collaborative scenarios. Interactive plans can potentially coordinate the efforts between multiple human users and an AI agent. Project documents are often highly collaborative artifacts. Interactive plans can play a similar role to existing task management tools like Trello [^11] https://trello.com/ that serve as coordinative artifacts (Bardram and Bossen, 2005; Schmidt and Wagner, 2002) and broaden common ground (Olson and Olson, 2000) in teams. They can help plan to-dos for the team, consolidate outputs in a central location, and display progress as items in the plan get completed.
Novel challenges arise for the agent in these collaborative scenarios. When proposing plans, the agent not only has to reason about whether a step should be assigned to a human user, but also who that will be. To do this, the agent may leverage context in the form of populated profiles of team members (e.g., researchers’ webpages and publication records, LinkedIn profiles outside of research). The agent may also infer expertise from the edit history of individuals within the document, similar to techniques for socially grounding AI chatbots in group chats using conversation history (Wang et al., 2024b). More generally, challenges in human-agent communication get exacerbated in collaborative settings. For example, the question of “which preferences should the agent respect?” (Bansal et al., 2024) is already challenging in a single-user setting because it requires the agent to elicit, preserve, and reason over task context and user preferences simultaneously. Shifting to collaborative settings necessitates keeping track of multiple sets of contexts and preferences and how they interact with each other. Collaborative use of interactive plans thus leave many open questions for future work.
@@ 10. Conclusion
In this paper, we presented Cocoa, an interactive document editing environment for scientific researchers to fluidly collaborate with an AI agent to tackle open questions and tasks within their research projects. Cocoa introduced a new interaction design pattern—interactive plans—that enabled Co-planning and Co-execution between a researcher and an AI agent.
Our lab ($n=16$) and field deployment ($n=7$) studies showed that participants not only engaged in collaborative planning and execution in Cocoa, but also interleaved co-planning and co-execution to progress on their tasks. Compared to a strong baseline with a more familiar chat interface, Cocoa enhanced agent steerability without sacrificing ease of use. In the lab study, participants’ main mode of co-execution was refining agent outputs, but when Cocoa was integrated into their research workflows in the field over a 7-day period, participants took advantage of user steps to guide the agent with their expertise and exert more of their agency. Overall, our work demonstrates the potential of interactive plans as a novel design pattern to flexibly mediate human and AI agency. Importantly, our work offers many practical takeaways for the design of agentic AI systems, such as leveraging user steps and advanced planning for agent cost management and oversight.
Agentic AI systems have exciting potential to transform our digital experiences and advance long-standing visions held by HCI and AI communities. However, these transformations and advancements also demand renewed attention to strategies for ensuring effective collaboration between human users and AI agents in the real world. Our work advances emerging efforts in this new frontier of human-AI collaboration.
@@ Appendix A Frequently Asked Questions
A.1. What’s the novelty of Cocoa compared to “deep research” systems?
It’s true that Cocoa and deep research systems can be used for similar tasks (e.g., literature review) and might even obtain similar results. However, deep research systems (at least at the time of writing) use fully automated AI workflows and do not accommodate user steering during the research process. Deep research systems may engage in some back and forth with users before the research process to clarify requests, or allow the user to follow up after the process ends, but does not afford user interaction during the process itself.
Why might this be undesirable? Let’s say you want to write a literature review on whether buttons with icons and labels are more usable than buttons without labels, or labels without icons. This is a real example that OpenAI has on their Deep Research release announcement webpage (OpenAI, 2025a). As an HCI researcher, you know of some reputable venues and authors you can rely on, but the agent used sources you might view as less credible (see sources in Deep Research’s response in the UX Design tab of (OpenAI, 2025a)). If Deep Research offered the affordances, you could guide the agent with your expertise during the research process to steer it towards using your trusted sources, which can also help improve the quality of its output. Without these affordances, you could follow up via prompting afterwards, but 1) there’s no guarantee that the agent will reliably follow those new instructions, and 2) re-running the research process is time- and resource-intensive. That’s why we draw on the computational notebook metaphor in Cocoa: interactive plans allow you to iteratively hone and refine the agent’s trajectory and outputs to bolster steerability, control, efficiency, and quality.
A.2. Can a user technically co-plan and co-execute with an agent via chat?
Technically, chat interfaces allow the user to ask the agent for a plan and prompt the agent to regenerate the plan before executing it (which has some aspects of co-planning). The user can also direct the agent to save some steps for themselves and prompt the agent to modify its outputs for steps assigned to the agent (which has some aspects of co-execution). However, these interactions are quite cumbersome in chat and not aligned with users’ mental models of chat interfaces—we did not witness any participants using our baseline system this way. Interactive plans are a fundamentally distinct design pattern designed to support interleaved co-planning and co-execution, and elicits different user behaviors than chat interfaces (see Figures 12 and 13 for evidence of this).
A.3. Do you see Cocoa’s lack of improvement over the baseline for general utility as a negative result?
No, we do not. General utility is a combination of three measures we obtained by asking participants the following Likert scale questions: 1) I’ve developed a better understanding of potential solutions to this particular problem, 2) I obtained new and useful insights from using system, and 3) The system’s outputs were useful for exploring the problem. It is unsurprising that there was no significant difference in answers to these questions, because these aspects are often bottlenecked by the performance of the underlying agent, and we used the same agent in both Cocoa and the baseline. However, we can improve aspects such as steerability with interface-level advances, which we have shown in this work. If general utility had gone up in Cocoa, it would have been a pleasant surprise, but it is understandable if it did not.
A.4. Why is it important to get the user involved? Can’t we just let AI automate everything?
@@
A.5. Why didn’t you use ChatGPT or a similar publicly available AI chatbot for your baseline?
Our baseline’s interface closely matches that of ChatGPT and other AI chatbots, but the underlying system is different. We use an LLM agent powered by GPT-4o, but specifically trained it to use the Semantic Scholar API as tools to reliably perform research-related tasks such as searches for papers/authors/topics, sorting papers by various metrics including citation count, summarizing papers, comparing papers, retrieving prominent papers by a particular author, etc. See Section 4.4 for technical details. However, our main contribution is not building a new LLM agent nor advancing agent capabilities—rather, it is to explore cost and benefit tradeoffs between interaction techniques for humans to work with AI agents. We use the same underlying agent in Cocoa, so we kept this factor constant across our system and baseline.